\vspace{.4in}
\begin{thebibliography}{99}
\bibitem{Tanaka2019} M. Tanaka, M. Takechi, A. Homma, M. Fukuda, 
D. Nishimura, T. Suzuki, Y. Tanaka, T. Moriguchi,
D.S. Ahn, A. Aimaganbetov, M. Amano, H. Arakawa, S. Bagchi, K.-H. Behr, N. Burtebayev,
K. Chikaato, H. Du, S. Ebata, T. Fujii, N. Fukuda, H. Geissel, T. Hori, W. Horiuchi, S. Hoshino,
R. Igosawa, A. Ikeda, N. Inabe, K. Inomata, K. Itahashi, T. Izumikawa, D. Kamioka, N. Kanda,
I. Kato, I. Kenzhina, Z. Korkulu, Y. Kuk, K. Kusaka, K. Matsuta, M. Mihara, E. Miyata, D. Nagae,
S. Nakamura, M. Nassurlla, K. Nishimuro, K. Nishizuka, K. Ohnishi, M. Ohtake, T. Ohtsubo, S. Omika,
H.J. Ong, A. Ozawa, A. Prochazka, H. Sakurai, C. Scheidenberger, Y. Shimizu, T. Sugihara,
T. Sumikama, H. Suzuki, S. Suzuki, H. Takeda, Y.K. Tanaka, I. Tanihata, T. Wada, K. Wakayama,
S. Yagi, T. Yamaguchi, R. Yanagihara, 
Y. Yanagisawa, K. Yoshida, and T.K. Zholdybayev, arXiv:1911.05262 [nucl-ex].
(to be published in Phys. Rev. Lett.)

\bibitem{Garcia2016} R. F. Garcia Ruiz, M. L. Bissell, K. Blaum, A. Ekstroem, N. Froemmgen, G. Hagen, M. Hammen,
K. Hebeler, J. D. Holt, G. R. Jansen, M. Kowalska, K. Kreim, W. Nazarewicz, R. Neugart,
G. Neyens, W. Nörtershäuser, T. Papenbrock, J. Papuga, A. Schwenk, J. Simonis,
K. A. Wendt and D. T. Yordanov, Nat. Phys. 12 (2016) 594.


\bibitem{Shalit1959} A. de-Shalit, Phys. Rev. 113 (1959) 547.

\bibitem{Abbas1980} A. Abbas and L. Zamick, Phys. Rev. C 21 (1980) 731. 

\bibitem{Zamick1977} L. Zamick, M. Golin and S. Moszkowski, Phys. Lett., B 66 (1977) 116.


\bibitem{Fujita1988} Y. Fujita, M. Fujiwara, S. Morinobu, T. Yamazaki, T. Itahashi, and H. Ikegami and 
S. I Hayakawa, Phys. Rev. C 37 (1988) 45.

\bibitem{nndc} www.nndc.bnl.gov/nudat2/

\bibitem{Burrows1993} W. Burrows, Nucl. Data Sheets 68 (1993) 1.



\end{thebibliography}

\documentclass[12pt]{article}
\usepackage{amssymb,amsmath}
\usepackage{graphicx}
\usepackage{ulem}
\usepackage{cite}

\begin{document} 
\begin{center} 
{\bf\Large { Extremely large matter radii in $^{49-51}$Ca isotopes and the $0^{+}$ breathing 
mode states  of $^{48}$Ca}} 
\end{center}
\vskip .1 cm 
\begin{center}
{\bf Syed Afsar Abbas}\\ 
Centre for Theoretical Physics, Jamia Millia Islamia University, New Delhi-110025, India\\
(email: drafsarabbas@gmail.com)\\
\vskip .1 cm
{\bf Anisul Ain Usmani, Usuf Rahaman}\\
Department of Physics, Aligarh Muslim University, Aligarh-202002, India
\end{center} 

\vspace{.05in} 
\begin{center} 
{\bf Abstract}
\end{center}
Through inelastic scattering cross section measurements for $^{42-51}$Ca on
a carbon target at 280 MeV/nucleon recently, Tanaka {\it et al.} [ arXiv:1911.05262 [nucl-ex]], 
in a very significant experiment, have demonstrated
large swelling of doubly magic $^{48}$Ca core in calcium isotopes beyond N=28. 
The matter radii observed in these experiments, are surprisingly much larger than the corresponding, already amazingly large 
charge radii of the same
calcium isotopes, by Garcia {\it et al.}
[Nat. Phys. 12 (2016) 594]. Here we propose a novel solution, wherein the breathing mode
states $0^{+}$ of $^{48}$Ca, provide a global and consistent solution of this matter radii conundrum.

\vskip .8 cm
\vspace{0.1in} 
{\bf PACS}: 21.60.-n, 21.10-k, 21.90.+f

\vspace{0.2in}
{\bf Keywords}: Charge radius, matter radius, calcium isotopes, 
inelastic scattering cross section, breathing mode, magicity,
E2 effective charge.

\newpage

Recently, Tanaka {\it et al.} ~\cite{Tanaka2019} have performed measurements of interaction 
cross section, $\sigma_{I}$  for $^{42-51}$Ca, on a neutral carbon target at 280 MeV/nucleon, 
in a groundgreaking first time experiment. Interaction cross section provides informations regarding RMS radius
of nucleon density distributions $<r^{2}_{m}>^{\frac{1}{2}}$, which is referred to as the
``matter radius'' here. This experiment complements the recent determination of charge
radii of the same calcium isotopes (plus $^{52}$Ca also), recently done by 
Garcia {\it et al.}~\cite{Garcia2016}. Surprisingly large charge radii were observed
in calcium isotopes $^{49}$Ca to $^{52}$Ca by Garcia {\it et al.}~\cite{Garcia2016}.
That has presented a great challenge to theoretical physics. Already a conundrum
as it was, in addition, Tanaka {\it et al.} ~\cite{Tanaka2019} have found that the matter radii
of the corresponding isotopes $^{49}$Ca to $^{51}$Ca, are amazingly, 
even much larger than the
corresponding charge radii. This presents a most challenging problem to nuclear 
structure physics.

Through meticulous calculations, Tanaka {\it et al.}~\cite{Tanaka2019} have 
demonstrated that these very large matter radii, cannot be taken as being due to one-,
two- and three-neutron haloes. They have also shown that this cannot be considered as
a deformation effect on these nuclei. What they have convincingly demonstrated
however is, that this is due to substantial swelling of the bare $^{48}$Ca 
core in these calcium isotopes.

So what can be the cause of this swelling of $^{48}$Ca core in these isotopes?
Getting motivation from the concept of E2 effective charge in nucleii, here
we propose a model wherein the breathing mode states $0^{+}$ provide a 
consistent understanding of this puzzling situation made explicit by 
Tanaka {\it et al.} ~\cite{Tanaka2019}.
One may ask, will this large enhancement in matter radii 
continue unabated as we add more and more neutrons to 
$^{48}$Ca core? And whether this will present us with a new calcium-radii-catastrophe?
Our model makes prediction that experimentalists shall find continuous large 
swelling of radii in $^{52}$Ca, $^{53}$Ca and $^{54}$Ca nuclei. 
After which there shall be 
no swelling. Hence, we shall show that our model is not only able to provide
a microscopic  understanding of what has already been demonstrated by 
Tanaka {\it et al.}, but provide new challenges for experimentalists 
to confirm in the future.

The idea of effective charge stems from a desire to explain electromagnetic 
transitions and moments totally within the framework of the shell model.
For example,  we renormalize the quadrupole operator
$\sum_{p}r^{2}Y_{2,0}$ to $(1+e_{p})\sum_{p}r^{2}Y_{2,0} + e_{n}\sum_{n}r^{2}Y_{2,0}$
where the effective charge correction for proton and neutron are 
$e_{p}$ and $e_{n}$ respectively. Note that the total effective charge of proton is $(1+e_{p})$.
It has been emperically determined that $e_{n} \approx +\frac{1}{2}$ and 
$e_{p} \approx +\frac{1}{2}$~\cite{Shalit1959,Abbas1980,Zamick1977}. One has to visualize, for
example, that in $^{17}$O, the $9^{th}$ valence neutron polarizes the $^{16}$O
core so strongly, that it gets highly deformed. Then we transfer this large deformation,
phenomenologicaly by treating the uncharged valence neutron 
to become a highly charged entity of magnitude 
$+\frac{1}{2}$, while still treating the core $^{16}$O as being a ground state spherical
nucleus~\cite{Shalit1959,Abbas1980,Zamick1977}.

In a similar manner, we treat here the $29^{th}$ valence neutron in 
$^{49}_{20}{\rm Ca}_{29}$, to sit outside the magical core nucleus
$^{48}_{20}{\rm Ca}_{28}$ and have an analogous strong deforming effect on it.
This "deforming'' effect, as ascertained by Tanaka {\it et al.}~\cite{Tanaka2019},
is not like that of  $^{16}$O in the case of effective charge  of $^{17}$O, but to 
``polarize'' it so that the core nucleus $^{48}_{20}{\rm Ca}_{28}$ 
``breathes'' out from the ground state 
$0^{+}$, to the first excited $0^{+}_{1}$ collective state of 
$^{48}_{20}{\rm Ca}_{28}$. As a consequence we treat the 
$29^{th}$ valence neutron to just sit smugly on the surface,
while the core has expanded out substantially to mock up the $0^{+}_{1}$ collective
state. This is its ground state now. Thus this is the source of large matter radius as observed 
by Tanaka {\it et al.} in $^{49}$Ca.

Next, in $^{50}_{20}{\rm Ca}_{30}$, now two neutrons sit smugly on the surface of the
core $^{48}_{20}{\rm Ca}_{28}$ nucleus, which then gets even more strongly ``polarized'' by these
(two neutrons), which then breathes out to the next $0^{+}_{2}$ collective state in 
$^{48}_{20}{\rm Ca}_{28}$, thereby providing a much larger matter radius. Similarly in 
$^{51}_{20}{\rm Ca}_{31}$, the valence neutrons sitting smugly on the surface of 
$^{48}_{20}{\rm Ca}_{28}$ would have induced still larger polarization of the core, to breath
out to the next collective $0^{+}_{3}$ state of $^{48}$Ca.

So having explained very large matter radii of $^{49}$Ca, $^{50}$Ca, $^{51}$Ca, we ask as to how long would this large expansion continue? That is,
will this expansion continue in uncontrolled manner in 
$^{52}$Ca, $^{53}$Ca, $^{54}$Ca,..., $^{59}$Ca, $^{60}$Ca? If so, we would have at hand, a 
genuine calcium-radii-catastrophe! One may be reminded of, the end $19^{th}$ century Ultraviolet Catastrophe of the black-body radiation.

As per the phenomenological model proposed here, the valence neutrons on top of N=28 in 
$^{48}_{20}{\rm Ca}_{28}$ would continue to polarize this core strongly, 
so as to go to higher and higher breathing mode states of
$^{48}_{20}{\rm Ca}_{28}$, as long as these states are available. As per data 
displayed below in Table 1, we give the $0^{+}_{n}$ excited breathing mode states of
$^{48}_{20}{\rm Ca}_{28}$, which have been determined experimentally
~\cite{Fujita1988,Burrows1993,nndc}.

\begin{table}
\centering
\caption{Experimental breathing mode states $0^{+}$ of $^{48}$Ca}
\vspace{.1in}
\renewcommand{\tabcolsep}{1.0cm}
\renewcommand{\arraystretch}{1.5}

\begin{tabular}{cc}
\hline\hline
$0^{+}_{1}$ & 4.284 MeV \\
$0^{+}_{2}$ & 5.461 MeV \\
$0^{+}_{3}$ & 11.945 MeV \\
$0^{+'}_{3}$ & 11.967 MeV \\
$0^{+'}_{4}$ & 12.318 MeV \\
$0^{+'}_{5}$ & 12.565 MeV \\
$0^{+'}_{6}$ & 12.869 MeV \\
\hline\hline
\label{tab1}
\end{tabular}
\end{table}

The breathing mode states $0^{+}_{n}$ of $^{48}$Ca have been taken 
from Ref.~\cite{Fujita1988,nndc}. Note the $0^{+}_{3}$ and $0^{+'}_{3}$ at 
11.945 MeV and 11.967 MeV respectively, appear as almost being degenerate. 
The role of these two as being one, is made 
sharper by the fact that in Burrows~\cite{Burrows1993},
$\sigma(\theta)$ in $(p,p^{'})$, $(\alpha,\alpha^{'})$ show oscillatory
pattern and are well fitted by DWBA by assuming and using $0^{+'}_{3}$, $0^{+'}_{4}$, 
$0^{+'}_{5}$, $0^{+'}_{6}$ states (while $0^{+}_{3}$ is not used). 
Note that it is possible that Giant Quadrupole
Resonance states and Giant Monopole Resonance states (breathing mode states) may mingle
with each other in a wide range of excitation energy in $^{48}$Ca ~\cite{Fujita1988}.
Thus if $0^{+}_{3}$ and $0^{+'}_{3}$ states are treated as one, 
the total number of breathing mode states
in $^{48}$Ca are six.

Thus the three observed large matter radii as obtained by Tanaka {\it et al.}~\cite{Tanaka2019}
in $^{49}$Ca, $^{50}$Ca, $^{51}$Ca, would be explained by the three breathing mode states
in $^{48}$Ca at $0^{+}_{1}$ = 4.284 MeV, $0^{+}_{2}$ = 5.461 MeV, 
$0^{+}_{3}$ = 11.945/11.967 MeV. Given the fact that we have three more states
$0^{+}_{4}$ = 12.318 MeV, $0^{+}_{5}$ = 12.565 MeV, $0^{+}_{6}$ = 12.869 MeV, our model
here predicts that one would observe still larger matter radii in 
$^{52}$Ca, $^{53}$Ca, $^{54}$Ca. Thus this expansion will stop at N=34, and
after that there shall be no further swelling, as per our phenomenological model prediction here.


\vskip .5 cm
\begin{center}
\end{center}

\end{document}